OpenStudy (anonymous):

find the locus of a point whose distance from (0,4) is 2/3 times the distance from the line y-9=0. name the curve

OpenStudy (anonymous):

step 1 let the moving point be (h,k)...

OpenStudy (anonymous):

step 2 (use given statements) distance of (h,k) from (0,4) which is sqrt((h-0)^2 +(k-4)^2) and distance of (h,k) from the line y-9=0 is |k-9|

OpenStudy (anonymous):

step3 condition(s)

OpenStudy (anonymous):

sqrt((h-0)^2 +(k-4)^2) =2/3 ( |k-9|)

OpenStudy (anonymous):

step4 change h to x and k to y and simplify further to get the required equation of the locus..

OpenStudy (anonymous):

sqrt((x-0)^2 +(y-4)^2) =2/3 ( |y-9|) squaring we have x^2 + (y-4)^2 =4/9 (y-9)^2...

OpenStudy (anonymous):

hope the steps are clear to you...

OpenStudy (anonymous):

thank you @matricked

OpenStudy (anonymous):

YW